The circumferences of two circles are 3 m. and 9πm. What is the ratio of the areas of the circle?


1
Expert's answer
2021-01-27T09:48:03-0500

If the circumferences of two circles are 3 m and 9πm, then their radiuses are 1.5/π and 4.5 m ("r = c/2\pi" ).

The ratio of the areas of the circle is:

"\frac{(1.5/π)^2}{4.5^2} = \frac{1}{9π^2}".